OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of two-thirds? 2x + 3y = 3 2x – 3y = 3 3x + 2y = 3 3x – 2y = 3

OpenStudy (anonymous):

\[Slope(m) = \frac{2}{3}\] Equation of line with one point given and a slope is given by: \[\large y - y_1 = m(x - x_1)\] So: \[y - 3 = \frac{2}{3}(x - 6)\] Multiply by 3 both the sides: \[3y - 9 = 2x - 12\] \[2x - 3y = 3\]